# Robust Weight Initialization for Tanh Neural Networks with Fixed Point Analysis

### v1

@@ 5 Conclusion
In this paper, we have introduced a novel weight initialization method for tanh FFNNs, grounded in the theoretical analysis of fixed points of the $\tanh(ax)$ function. Through our fixed-point analysis, we established conditions under which the vanishing or exploding of activation values can be prevented, even as the depth of the network increases.
Our proposed method exhibits strong robustness to variations in network size, as demonstrated across a variety of FFNN configurations and benchmark datasets, including MNIST, Fashion MNIST, and CIFAR-10. In contrast to Xavier initialization, which struggles to maintain stable performance as network depth increases, the proposed method consistently achieves superior results by preserving activation values. Furthermore, we explored the impact of the initialization hyperparameter $\sigma_{z}$ on the distribution of activation values. We demonstrated both theoretically and experimentally that the choice of $\sigma_{z}$ plays a significant role in maintaining the proper range of activations, balancing between vanishing and saturation. In the context of PINNs, the proposed initialization method shows improved performance in solving PDEs such as the Burgers’ equation and the Allen-Cahn equation. By maintaining a stable loss function and achieving faster convergence compared to Xavier initialization, our method demonstrates its practical utility in training networks for physical systems.
A key advantage of the proposed method lies in its robustness to network depth and width, significantly reducing the need for extensive hyperparameter tuning. By maintaining stable performance across varying network configurations, our approach helps to minimize the time and effort spent on searching for optimal network architectures, allowing researchers to focus on model design and other aspects of the training process. This makes the proposed method particularly valuable in large-scale and resource-constrained applications where efficient training is critical.
Appendix A Proofs of the theoretical results
@@ A.1 Proof of Lemma 1
We define $g(x)=\tanh(ax)-x$. Since $g(x)$ is continuous, and $g(-M)>0$, $g(M)<0$ for a large real number $M\in\mathbb{R}^{+}$, the Intermediate Value Theorem guarantees the existence of a point $x$ such that $g(x)=0$.
First, consider the case $0<a\leq 1$. Since $0<a\leq 1$, the derivative $g^{\prime}(x)=a\cdot\text{sech}^{2}(ax)-1$ satisfies $-1\leq g^{\prime}(x)\leq a-1<0$ for all $x$. Hence, $g(x)$ is strictly decreasing and therefore $g(x)$ has the unique root. At $x=0$, $\phi(0)=\tanh(a\cdot 0)=0$. Hence, $x=0$ is the unique fixed point.
Let us consider the case $a>1$. For $0<x\ll 1$, $\tanh(ax)-x\approx(a-1)x$. Since $a>1,\tanh(ax)-x>0$. On the other hand, since $|\tanh(ax)|<1$ for all $x$,
@@ A.2 Proof of Lemma 2
(2) Let $x_{0}<\xi_{a}$. Since $\phi^{\prime}(x)$ decreasing for $x\geq 0$, with $\phi^{\prime}(0)>1$ and $\xi_{a}$ is the unique fixed point for $x>0$, it holds that $x_{n}<x_{n+1}<\xi_{a}$ for all $n\in\mathbb{N}$. Thus, by the monotone convergence theorem, the sequence converges to the fixed point $\xi_{a}$. The proof is similar when $x_{0}>\xi_{a}$. ∎
Appendix B Additional experimental results
B.1 Preventing activation saturation via appropriate $\sigma_{z}$ tuning
@@
Figure 4: Here, ’STD’ refers to $\sigma_{z}$. (a) shows the PINN loss for the Burgers’ equation, using an FFNN with 30 layers and 32 nodes in each hidden layer. (b) shows the PINN loss for an FFNN with 30 layers, where the hidden layers alternate between 64 and 32 nodes, repeated 15 times. Each experiment was repeated 10 times with different random seeds.
B.2 Width independence in Classification tasks
Figure 5: Validation accuracy and loss are presented for tanh FFNNs with varying numbers of nodes $(2,8,32,128)$, each with 20 hidden layers. All models were trained for 20 epochs on the MNIST dataset, with 10 different random seeds.
Figure 6: Validation accuracy and loss are presented for tanh FFNNs with varying numbers of nodes $(2,8,32,128)$, each with 20 hidden layers. All models were trained for 20 epochs on the Fashion MNIST dataset, with 10 different random seeds.
Figure 7: Validation accuracy and loss are presented for tanh FFNNs with varying numbers of nodes $(2,8,32,128)$, each with 20 hidden layers. All models were trained for 20 epochs on the CIFAR-10 dataset, with 10 different random seeds.

### v2

@@ 5 Conclusion
In this study, we proposed a novel weight initialization method for tanh neural networks, baesed on a theoretical analysis of fixed points of $\tanh(ax)$ function. The proposed method is experimentally demonstrated to achieve robustness to variations in network size without normalization methods and to exhibit improved data efficiency. Therefore, the proposed weight initialization method reduces the time and effort required for training neural networks.
Acknowledgments
This work of Hyunwoo Lee and Hayoung Choi was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2022R1A5A1033624 and RS-2024-00342939). The work of Hyunju Kim was supported by the funds of the Open R&\&&D program of Korea Electric Power Corporation.(No-R23XO03).
Appendix A Analysis of Signal Propagation
In this section, we provide proofs for the statements presented in Section 3.1. Each proof follows from the fundamental properties of $\tanh(ax)$ and analytical derivations. An empirical analysis is performed on the activation distribution for normally and beta-distributed input data.
@@ A.1 Proof of Lemma 1
Define $g(x)=\tanh(ax)-x$. Since $g(x)$ is continuous, and $g(-M)>0$, $g(M)<0$ for a large real number $M$, the Intermediate Value Theorem guarantees the existence of a point $x$ such that $g(x)=0$.
First, consider the case $0<a\leq 1$. Since $0<a\leq 1$, the derivative $g^{\prime}(x)=a\cdot\text{sech}^{2}(ax)-1$ satisfies $-1\leq g^{\prime}(x)\leq a-1<0$ for all $x$. Hence, $g(x)$ is strictly decreasing and therefore $g(x)$ has a unique root. At $x=0$, $\phi(0)=\tanh(a\cdot 0)=0$. Hence, $x=0$ is the unique fixed point.
We consider the case $a>1$. For $0<x\ll 1$, $\tanh(ax)-x\approx(a-1)x$. Since $a>1,\tanh(ax)-x>0$. On the other hand, since $|\tanh(ax)|<1$ for all $x$,
@@ A.2 Proof of Lemma 2
(2) Let $x_{0}<\xi_{a}$. Since $\phi^{\prime}(x)$ is decreasing for $x\geq 0$, with $\phi^{\prime}(0)>1$ and $\xi_{a}$ is the unique fixed point for $x>0$, it holds that $x_{n}<x_{n+1}<\xi_{a}$ for all $n\in\mathbb{N}$. Thus, by the monotone convergence theorem, the sequence converges to the fixed point $\xi_{a}$. The proof is similar when $x_{0}>\xi_{a}$. By the monotone convergence theorem, the sequence also converges to the fixed point $\xi_{a}$. ∎
A.3 Proof of Proposition 3
Proof.
Then the inequality $\hat{\Phi}_{m}\leq\Phi_{m}\leq\tilde{\Phi}_{m}$ holds for all $m$. By Lemma 1, for any $x\geq 0$, we have $\lim_{m\to\infty}\hat{\Phi}_{m}=0$ and $\lim_{m\to\infty}\tilde{\Phi}_{m}=0$. Therefore, the Squeeze Theorem guarantees that $\lim_{m\to\infty}\Phi_{m}(x)=0$. ∎
A.4 Proof of Corollary 4
Proof.
A.5 Activation Distribution Based on Input Data Distribution
Figure 8: The activation values in the 1000th layer, with $32$ nodes per hidden layer, were analyzed using the proposed weight initialization method with $\sigma_{z}$ values of $0.0003$, $0.015$, $0.3$, and $3$. The upper row shows results for $3000$ input samples drawn from a standard normal distribution, while the lower row presents results for samples drawn from a Beta distribution with parameters $a=2.0$ and $b=5.0$.
Appendix B Classification Tasks
In this section, we present experimental results for the benchmark classification datasets MNIST, Fashion-MNIST, CIFAR-10, and CIFAR-100 when employing the proposed weight initialization method.
B.1 Width independence in Classification tasks
[…]
B.2 Non-uniform Hidden Layer Dimensions
Tanh neural networks have been less widely employed compared to ReLU networks due to their higher computational complexity, susceptibility to the vanishing gradient problem, and the superior empirical performance of ReLU in various deep learning tasks. However, the recent success of PINNs using tanh neural networks has renewed interest in their application. This section compares the performance of four initialization methods on architectures that are generally challenging to train: (1) tanh activation with Xavier initialization, (2) tanh activation with the proposed initialization, (3) ReLU activation with He initialization + BN, and (4) ReLU activation with orthogonal initialization. The experiments are performed on FFNN and autoencoder models.
FFNN. The experiment is performed using an FFNN with a structure in which hidden layers alternated between 16 and 4 nodes, repeated 50 times, and trained over 100 epochs. The results are shown in Figure 13 (a). Both Xavier and the proposed method successfully train the network, with the proposed method showing overall better performance. Due to its strong performance despite substantial variations in hidden layer sizes, the proposed method is further evaluated on autoencoders with large variations in layer sizes, as shown in Figure 13 (b).
Autoencoder. The autoencoder architecture consists of an encoder and a decoder, both employing batch normalization and dropout (0.2) for regularization. The encoder compresses the input through layers of 512, 256, and 128 units before mapping to a 64-dimensional latent space. The decoder reconstructs the input by symmetrically expanding the latent space through layers of 128, 256, and 512 units, followed by an output layer with sigmoid activation. In Figure 13 (b), the model is trained on the MNIST dataset with a batch size of $256$, while in (c), it is trained on the FMNIST dataset with a batch size of $512$.
(a) FFNN (MNIST)
(b) Autoencoder (MNIST)
(c) Autoencoder (FMNIST)
Figure 13: (a) Validation loss for an FFNN with alternating hidden layers of $16$ and $4$ nodes, repeated $50$ times, comparing four methods: Tanh with Xavier initialization, Tanh with the proposed initialization, ReLU with He initialization + BN, and ReLU with orthogonal initialization. (b) Validation loss for an autoencoder with encoder-decoder layers of $512$, $256$, $128$, and $64$ units, comparing the same four methods. (c) Same as (b), but on the FMNIST dataset.
Appendix C Physics-Informed Neural Networks
In this section, we present additional experiments employing the proposed weight initialization method in Physics-Informed Neural Networks (PINNs). Specifically, we empirically analyze its relationship with different activation functions (Swish, ELU, Sigmoid, and ReLU) and its impact on various PDE problems (Burgers’ equation, Allen-Cahn equation, Poisson equation, and Diffusion equation).
C.1 PDE Details
Here we present the differential equations that we study in experiments.
Allen-Cahn Equation. The diffusion coefficient is set to $d=0.01$. The initial condition is defined as $u(x,0)=x^{2}\cos(\pi x)$ for $x\in[-1,1]$, with boundary conditions $u(-1,t)=-1$ and $u(1,t)=-1$, applied over the time interval $t\in[0,1]$. The Allen-Cahn equation is expressed as
$\frac{\partial u}{\partial t}-d\frac{\partial^{2}u}{\partial x^{2}}=-\frac{u^{3}+u}{d},$
Burgers’ Equation. The Burgers’ equation, a viscosity coefficient of $\nu=0.01$ is employed. The initial condition is given by $u(x,0)=-\sin(\pi x)$ for $x\in[-1,1]$, with boundary conditions $u(-1,t)=0$ and $u(1,t)=0$ imposed for $t\in[0,1]$. The Burgers’ equation is expressed as
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}},$
Diffusion Equation. The diffusion equation includes a time-dependent source term and is defined over the spatial domain $x\in[-1,1]$ and temporal interval $t\in[0,1]$. The initial condition is specified as $u(x,0)=\sin(\pi x)$, with Dirichlet boundary conditions $u(-1,t)=0$ and $u(1,t)=0$. The diffusion equation is expressed as
$\frac{\partial u}{\partial t}-\frac{\partial^{2}u}{\partial x^{2}}=e^{-t}\left(\sin(\pi x)-\pi^{2}\sin(\pi x)\right),$
where $u(x,t)$ is the solution.
Poisson Equation. The Poisson equation is defined over the spatial domain $x\in[0,1]$ and $y\in[0,1]$. The Poisson equation is expressed as
$\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}=f(x,y),$
where $u(x,y)$ is the solution, and $f(x,y)$ is the source term given by
$f(x,y)=2\pi^{2}\sin(\pi x)\sin(\pi y).$
C.2 Effect of activation function on PINNs
We experimentally demonstrate that when employing the proposed weight initialization method, the absolute error between the exact solution and the PINN-predicted solution is smaller when using the tanh activation compared to ReLU, sigmoid, Swish (Ramachandran et al., 2017), and ELU (Clevert, 2015), as shown in Figure 14. The FFNN consists of 30 hidden layers (32 nodes each) and is trained for 300 iterations using Adam, followed by 300 iterations using L-BFGS. For the Burgers’ and diffusion equations, the PINN with tanh activation achieves the closest approximation to the exact solution. It is well established that PINNs with tanh activation effectively approximate solutions to the Burgers’ and diffusion equations. Given that the choice of activation function and weight initialization are dependent, this result provides valuable insights into the interaction between initialization methods and activation functions in PINNs..
Figure 14: Absolute error for (upper row) the Burgers’ equation and (lower row) the diffusion equation with varying activation functions. Values outside the color bar range are depicted in white.
Figure 15 presents experiments that employ the proposed initialization and Xavier initialization for PDE problems using the Swish activation function, which is widely used in PINNs.
(a) Diffusion equation
(b) Poisson equation
(c) Burgers’ equation
(d) Allen-Cahn equation
Figure 15: PINN loss for a Swish FFNN with (a) $20$ hidden layers, each containing $32$ nodes, and (b), (c) $3$ hidden layers, each containing $3$ nodes, and (d) $10$ hidden layers, each containing $32$ nodes.
C.3 $\sigma_{z}$ for Burgers’ Equation
@@
Figure 16: Here, STD refers to $\sigma_{z}$. (a) shows the PINN loss for the Burgers’ equation, using an FFNN with 30 layers and 32 nodes in each hidden layer. (b) shows the PINN loss for an FFNN with 30 layers, where the hidden layers alternate between 64 and 32 nodes, repeated 15 times.
Here, we present the PINN loss for the Burgers’ equation as a function of the initialization hyperparameter $\sigma_{z}$. As shown in Figure 16, the experimental results demonstrate that $\sigma_{z}$ significantly impacts training. Therefore, selecting an appropriate $\sigma_{z}$ is crucial for optimizing performance.
C.4 Absolute error for Burgers’ equation
Figure 17 demonstrates that increasing the number of collocation points allows PINNs with the proposed initialization to predict solutions with lower absolute error compared to those using Xavier initialization. The proposed method is experimentally demonstrated to achieve faster convergence speed and higher data efficiency.
Figure 17: Absolute error between the exact solution and the PINN-predicted solution for the Burgers’ equation with varying numbers of collocation points (3000, 10000, 20000, 50000) using (upper row) Xavier and (lower row) the proposed initialization. The FFNN has 30 hidden layers ($32$ nodes each) and is trained for $300$ iterations using Adam followed by $300$ iterations using L-BFGS.
Appendix D Examples of the matrix $\mathbf{D}^{\ell}$
An example of $\mathbf{D}^{\ell}$ from the initialization methodology proposed in Section 3.2 is presented. In the heatmap of Figure 18, white represents the value 00, while black represents the value $1$.
Figure 18: Examples of the matrix $\mathbf{D}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$ in Section 3.2 with $N_{\ell}=4,N_{\ell-1}=8$ (left), $N_{\ell}=4,N_{\ell-1}=4$ (middle), and $N_{\ell}=8,N_{\ell-1}=4$ (right), respectively.